OpenStudy (anonymous):

ou find the vehicle you are purchasing and its price. Vehicle: Chevy Volt Price: $39,145 Current interest rate: 3% Being a smart financial planner, you want to figure out how many months it will be until your principal is paid down to $10,000.00. Solve for t and show all of your work. Note that t will be negative because the number of months will decrease the principal.

OpenStudy (amistre64):

without making any payments, you can decrease it. what is your payment amount?

OpenStudy (anonymous):

A(t)= P (1+r/n)^nt my function is A(t)= 39,145(1+0.03/12)^12t Now P does not represent what I put down, it represents the cost of the vehical